two coils connected in parallel across 100 V supply take 10 A from the line. The power dissipated in one coil is 600 W.The resistance of the other is...?

Grade:10

1 Answers

Vikas TU
7 years ago
let current in coils be from krichff’s rule,
I and 10 – I
then,
600 = V^2/R
R = 10000/600 = > 100/6 = > 50/3 ohms.
also,
100 = I*(50/3)
I = 6 A in one coil and 4 A in other coil.
hence,
100 = 4*R
R = 25 ohms.